# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

### v2

@@ III-C Frustum-Range Operators
Data augmentation enables the model to learn more robust representations. Common augmentations in LiDAR segmentation typically operate on the global view, encompassing rotation, flipping, scaling, jittering, etc. In this work, we introduce two novel augmentations based on our proposed frustum-range view representation.
FrustumMix. Previous mixing strategies [44, 45] involve swapping two LiDAR scans at the point level, which may disrupt the internal structure within the frustum region. In this work, we propose FrustumMix, a method to swap frustum regions from two scenes. Specifically, given two LiDAR point clouds $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, we randomly split the point clouds into $M$ non-overlapping frustum regions along the inclination or azimuth direction via Eq. 1, denoted as $\mathcal{P}_{1}=\{\mathcal{A}_{1}^{1},...,\mathcal{A}_{1}^{M}\}$ and $\mathcal{P}_{2}=\{\mathcal{A}_{2}^{1},...,\mathcal{A}_{2}^{M}\}$. We then generate a new LiDAR scan $\hat{\mathcal{P}}$ by alternately swapping frustum regions from $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, as follows: $\hat{\widetilde{P}}=\mathcal{A}_{1}^{1}\cup\mathcal{A}_{2}^{2}\cup\mathcal{A}_{1}^{3}\cup\mathcal{A}_{2}^{4}\cup\cdots$. FrustumMix enhances context-awareness between frustums while preserving the invariant structure within the frustum.
@@
PTv2 [66] NeurIPS’22 Raw Points $\bullet$ $12.8$ M $4.7$ $\mathbf{70.3}$ $\mathbf{72.6}$ $\mathbf{80.2}$ $\mathbf{82.6}$
WaffleIron [11] ICCV’23 Raw Points $\bullet$ $6.8$ M $\mathbf{7.4}$ $68.0$ $70.8$ $79.1$
@@
2DPASS [67] ECCV’22 Multi-View $\bullet$ $26.5$ M $\mathbf{8.4}$ $69.3$ $72.2$ $\mathbf{79.4}$ $80.8$
@@
UniSeg [23] ICCV’23 Multi-View $\bullet$ $147.6$ M $6.9$ $\mathbf{71.3}$ $\mathbf{75.2}$ $\mathbf{83.5}$
@@
RangeViT [31] CVPR’23 Range View $\bullet$ $23.7$ M $10.0$ $60.7$ $64.0$ $75.2$ $53.6$ $66.5$
RangeFormer [32] ICCV’23 Range View $\bullet$ $24.3$ M $6.2$ $67.6$ $\mathbf{73.3}$ $78.1$ $80.1$ $63.0$
@@
RangeInterpolation. Previous range-view methods [58, 55, 32] often encounter a large number of empty pixels due to the sparsity of LiDAR points, resulting in noisy information. To mitigate this problem, we introduce RangeInterpolation, a technique that utilizes the range image to reconstruct surfaces in the LiDAR scan. Specifically, the point cloud $\mathcal{P}$ is first projected onto the range image $\mathcal{R}\in\mathbb{R}^{H\times W\times(3+L)}$. For empty pixel position $(u,v)$ in the range image $\mathcal{R}$, we establish a window with size $m\times n$ centered at $(u,v)$, where $m$ and $n$ are odd. Our aim is to interpolate the point falling into $(u,v)$ using the surrounding range information within the predefined window. Considering that not all pixels in the window contain valid values, we aggregate the information from non-empty pixels within the window for point interpolation using an average operation. During the training procedure, the semantic label of the interpolated point is determined via threshold-guided voting within the window. Specifically, we identify the most frequent category within the window and its frequency. If the frequency is below a threshold, we classify the interpolated point as located at the boundary and assign the label as ignored, which will not contribute to loss calculation. Otherwise, the label is determined by the category with the highest frequency.
@@ IV Experiments
FrustumMix Ours $\mathbf{55.8}$ $\mathbf{64.8}$ $\mathbf{65.2}$ $\mathbf{65.4}$ $\mathbf{61.2}$ $\mathbf{72.2}$ $\mathbf{74.6}$ $\mathbf{75.4}$ $\mathbf{46.6}$ $57.0$ $59.5$ $\mathbf{61.2}$
@@ IV-A Experimental Settings
Datasets. We conduct a comprehensive experiment on four popular LiDAR segmentation datasets. SemanticKITTI [28] is a large-scale outdoor dataset that comprises $22$ sequences collected from various scenes in Karlsruhe, Germany, using a Velodyne HDL-64E LiDAR sensor. Officially, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) for validation, and sequences $11$ to $21$ ($20,351$ scans) for online testing. The original dataset is annotated with $28$ classes, though only $19$ merged classes are used for single-scan evaluation. The vertical FOV is $[-25^{\circ},3^{\circ}]$. nuScenes [43] is a multimodal dataset that is widely used in autonomous driving scenarios. It contains LiDAR point clouds captured around streets in Boston and Singapore, provided by 32-beam LiDAR sensors. The dataset includes $1,000$ driving scenes with sparser points. It is annotated with $32$ classes and only $16$ semantic categories are used for evaluation. The vertical FOV is $[-30^{\circ},10^{\circ}]$. SemanticPOSS [64] is a more challenging benchmark collected at Peking University and features much smaller and sparser point clouds. It consists of $2,988$ LiDAR scenes with numerous dynamic instances. Officially, it is divided into $6$ sequences, with sequence $2$ used for evaluation with $13$ merged categories. The vertical FOV is $[-16^{\circ},7^{\circ}]$. ScribbleKITTI [63] shares the same scenes with SemanticKITTI [28] but only provides weak annotations with line scribbles. It contains $189$ million labeled points, with approximately $8.06\%$ labeled points available during training. Additionally, we conduct a comprehensive out-of-distribution robustness evaluation experiment on SemanticKITTI-C and nuScenes-C from the Robo3D [71] benchmark. Each dataset contains eight corruption types, including “fog”, “wet ground”, “snow”, “motion blur”, “beam missing”, “crosstalk”, “incomplete echo”, and “cross sensor”.
@@
Implementation Details. We implement FRNet using the popular MMDetection3D [72] codebase and employ CENet [58] as the 2D backbone to extract frustum features. The 2D representation shape is set to $64\times 512$ for SemanticKITTI [28] and ScribbelKITTI [63], and $32\times 480$ for nuScenes [43] and SemanticPOSS [64], respectively. For optimization, We choose AdamW [73] as our default optimizer with an initial learning rate of $0.01$. We utilized the OneCycle scheduler [74] to dynamically adjust the learning rate during training. All models are trained using four GPUs for $50$ epochs on SemanticKITTI [28] and ScribbelKITTI [63], and $80$ epochs on nuScenes [43] and SemanticPOSS [64], respectively. The batch size is set to $4$ for each GPU. Furthermore, we implement a faster variant, dubbed Fast-FRNet, which features a smaller 2D representation resolution and scaled-down 2D backbone. Specifically, the size of the 2D representation is adjusted to $32\times 360$ for all datasets, and the 2D backbone is a variant of ResNet18 [75], following the design of CENet [58].
@@ IV-B Comparative Study
Comparisons to State of the Arts. We begin by comparing FRNet with state-of-the-art LiDAR segmentation methods across different representations, including sparse-voxel, point-view, multi-view, and range-view representations, on both fully-supervised and weakly-supervised benchmarks. Tab. I provides a summary of all the results. We evaluate the Frame Per Second (FPS) on a single GeForce RTX 2080Ti GPU for fair comparison. For SemanticKITTI [28] and nuScenes [43], we report mIoU scores on both the val and test sets. For ScribbleKITTI [63] and SemanticPOSS [64], which share the same data on val and test sets, we report both mIoU and mAcc scores on the val set. Compared with multi-view methods, although FRNet does not exhibit a significant advantage in performance, its parameters and FPS are notably superior. In comparison with sparse-voxel-view and range-view methods, FRNet achieves appealing performance while still maintaining high efficiency. Additionally, Fast-FRNet demonstrates the fastest inference speed with only a minor drop in performance.
@@
Label-Efficient LiDAR Semantic Segmentation. Recently, semi-supervised learning has been explored in LiDAR segmentation. In this work, we adopt the approach of LaserMix [44] to apply FRNet to the semi-supervised segmentation task, as shown in Tab. II. Firstly, we utilize Cylinder3D [16] as the backbone and leverage the proposed FrustumMix technique to mix labeled and unlabeled data for consistency learning. Notably, FrustumMix demonstrates superior performance compared to the LaserMix [44] method. Subsequently, we employ FRNet as the backbone and assess various strategies for mixing labeled and unlabeled scans, including PolarMix [45], LaserMix [44], and our FrustumMix. The Sup.-only baseline results are obtained without FrustumMix and RangeInterpolation as data augmentation techniques. Results show that FrustumMix consistently achieves the best scores across all data splits, particularly with limited annotations.
Out-of-Distribution Robustness. Robustness, which reflects a model’s ability to generalize under different conditions, is crucial for automotive security systems, particularly in extreme weather conditions [76, 77]. To evaluate the out-of-distribution robustness of FRNet, we utilize the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [71]. The models trained on SemanticKITTI [28] and nuScenes [43] are directly applied to evaluate the performance on SemanticKITTI-C and nuScenes-C [71], respectively. As shown in Tab. III, FRNet achieves promising performance across most corruption types and demonstrates superiority over recent works employing various LiDAR representations, including sparse voxel [17, 15, 16], range view [55, 58], multi-view fusion [67], and raw points [11].
@@
[…]
@@
IV-C Ablation Study
Component Designs. Tab. IV summarizes the ablation results of each component in the FRNet architecture. Firstly, We directly apply the frustum feature encoder to CENet [58], incorporating KNN post-processing to predict semantic labels over the entire point cloud, yielding $62.7\%$ and $72.8\%$ mIoU on SemanticKITTI [28] and nuScenes [43], respectively. Subsequently, by integrating the point-frustum fusion to hierarchically update point features, the semantic results are predicted with the updated point features in an end-to-end manner, resulting in an improvement of $1.6\%$ and $0.8\%$ mIoU, respectively. To regularize the frustum features, the frustum-level supervision is introduced and it leads to significant gains of $1.9\%$ and $2.8\%$ mIoU, respectively. The fusion head module, which fuses multiple features at different levels for more accurate prediction, brings an additional $0.8\%$ mIoU gain on both datasets. Furthermore, the proposed RangeInterpolation, aimed at alleviating empty pixels in the 2D representation by reconstructing semantic surfaces via surrounding range information, provides a performance boost of approximately $0.5\%$ mIoU. Finally, adopting Test Time Augmentation during inference, following prior works, brings an improvement of $1.1\%$ and $1.3\%$ mIoU, respectively.
@@ IV-C Ablation Study
TABLE V: Ablation study on window settings in RangeInterpolation on the val set of SemanticKITTI [28]. Green pixel denotes the empty pixel needed to be interpolated. Violet pixels are the surrounding pixels used to generate the point. All mIoU scores are given in percentage (%).
@@
RangeInterpolation Window Settings. RangeInterpolation reconstructs semantic surfaces based on the range image, facilitating a compact 2D representation for coherent semantic learning. Tab. V showcases the performance of RangeInterpolation with different window settings. Notably, We observe consistent performance improvements when considering interpolation along the horizontal or vertical direction exclusively. Specifically, configurations employing horizontal windows, such as the $1\times 3$ and $1\times 5$ settings, demonstrate superior results compared to those focusing on the vertical direction ($3\times 1$ and $5\times 1$ windows). We conjecture that points aligned along the same laser beams often share similar attributes, leading to more accurate and coherent semantic surfaces. Conversely, larger window sizes, such as the $3\times 5$, $3\times 3$, and $5\times 5$ configurations, tend to exhibit non-uniform point distributions within the window, resulting in noisy points that fail to accurately represent the expected region and consequently leading to a drop in performance.

### v3

@@ III-C Frustum-Range Operators
Data augmentation enables the model to learn more robust representations. Common augmentations in LiDAR segmentation typically operate from a global perspective, including rotation, flipping, scaling, jittering, etc. In this work, we introduce two novel augmentations based on our proposed frustum-range view representation.
FrustumMix. Previous mixing strategies [48, 49] perform point-level swapping between two LiDAR scans. While these methods achieve notable improvements in various LiDAR representations, they are not well aligned with the frustum representation, potentially causing two distinct scenes to mix within the same frustum region. This misalignment disrupts both the semantic coherence and geometric structure within the frustum region. To this end, we propose FrustumMix, a fine-grained extension of LaserMix [48] that aligns better with the frustum representation. Specifically, given two LiDAR point clouds $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, we randomly split the point clouds into $M$ non-overlapping frustum regions along the inclination or azimuth direction, denoted as $\mathcal{P}_{1}=\{\mathcal{A}_{1}^{1},...,\mathcal{A}_{1}^{M}\}$ and $\mathcal{P}_{2}=\{\mathcal{A}_{2}^{1},...,\mathcal{A}_{2}^{M}\}$. We then generate a new LiDAR scan $\hat{\mathcal{P}}$ by alternately swapping frustum regions from $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, as follows: $\hat{\mathcal{P}}=\mathcal{A}_{1}^{1}\cup\mathcal{A}_{2}^{2}\cup\mathcal{A}_{1}^{3}\cup\mathcal{A}_{2}^{4}\cup\cdots$. This approach ensures that mixing occurs at the level of frustum units, preserving semantic coherence and structural details entirely within each frustum region, thereby maintaining alignment with the frustum representation. Fig. 5 provides an illustration of FrustumMix. By enhancing context-awareness between frustums while preserving intra-frustum structural invariance, FrustumMix achieves improved segmentation performance and better representation learning.
@@
PTv2 [70] NeurIPS’22 Raw Points $\bullet$ $12.8$ M $4.7$ $70.3$ $72.6$ $80.2$ $82.6$
WaffleIron [11] ICCV’23 Raw Points $\bullet$ $6.8$ M $7.4$ $68.0$ $70.8$ $79.1$
PTv3 [12] CVPR’24 Raw Points $\bullet$ $46.2$ M $\mathbf{20.2}$ $\mathbf{70.8}$ $\mathbf{74.2}$ $\mathbf{80.4}$ $\mathbf{82.7}$
@@
2DPASS [71] ECCV’22 Multi-View $\bullet$ $26.5$ M $8.4$ $69.3$ $72.2$ $\mathbf{79.4}$ $80.8$
@@
UniSeg [24] ICCV’23 Multi-View $\bullet$ $147.6$ M $6.9$ $71.3$ $75.2$ $83.5$
TASeg [73] CVPR’24 Multi-View $\bullet$ $46.7$ M $6.1$ $\mathbf{72.7}$ $\mathbf{76.5}$ $\mathbf{84.6}$
@@
[…]
@@
RangeInterpolation. Previous range-view methods [61, 59, 36] often encounter a large number of empty pixels due to the sparsity of LiDAR points, resulting in noisy information. To mitigate this problem, we introduce RangeInterpolation, a technique that utilizes the range image to reconstruct surfaces in the LiDAR scan. Specifically, the point cloud $\mathcal{P}$ is first projected onto the range image $\mathcal{R}\in\mathbb{R}^{H\times W\times(3+L)}$. For empty pixel positions $(u,v)$ in the range image $\mathcal{R}$, we establish a window of size $m\times n$ centered at $(u,v)$, where $m$ and $n$ are odd. We aim to interpolate the point falling into $(u,v)$ using the surrounding range information within the pre-defined window. Considering that not all pixels in the window contain valid values, we aggregate information from non-empty pixels within the window for point interpolation using an average operation. During the training procedure, the semantic label of the interpolated point is determined via threshold-guided voting within the window. Specifically, we identify the most frequent category within the window and its frequency. If the frequency is below a threshold, we classify the interpolated point as located at the boundary and assign the label as ignored, which will not contribute to loss calculation. Otherwise, the label is determined by the category with the highest frequency. As shown in Fig. 6, RangeInterpolation produces coherent semantic information in the range image, resulting in a more compact representation.
@@ IV Experiments
Sup.-only - PTv3 [12] $42.7$ $60.9$ $62.6$ $64.1$ $48.0$ $67.8$ $71.5$ $75.9$ $40.7$ $54.6$ $56.2$ $58.9$
PolarMix [49] NeurIPS’22 $47.2$ $61.6$ $63.1$ $64.8$ $53.4$ $69.3$ $72.0$ $76.1$ $42.5$ $55.8$ $56.9$ $59.6$
LaserMix [48] CVPR’23 $51.7$ $63.3$ $64.8$ $65.3$ $57.4$ $71.8$ $74.2$ $77.4$ $45.0$ $56.9$ $58.3$ $60.7$
FrustumMix Ours $54.2$ $64.7$ $\mathbf{65.6}$ $\mathbf{66.1}$ $60.3$ $\mathbf{72.3}$ $\mathbf{75.6}$ $\mathbf{77.7}$ $46.2$ $57.5$ $60.0$ $\mathbf{61.6}$
@@
FrustumMix Ours $\mathbf{55.8}$ $\mathbf{64.8}$ $65.2$ $65.4$ $\mathbf{61.2}$ $72.2$ $74.6$ $75.4$ $\mathbf{46.6}$ $57.0$ $59.5$ $61.2$
@@ IV-A Experimental Settings
Datasets. We conduct comprehensive experiments on four popular LiDAR segmentation datasets. SemanticKITTI [32] is a large-scale outdoor dataset that comprises $22$ sequences collected from various scenes in Karlsruhe, Germany, using a Velodyne HDL-64E LiDAR sensor. Officially, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) for validation, and sequences $11$ to $21$ ($20,351$ scans) for online testing. The original dataset is annotated with $28$ classes, though only $19$ merged classes are used for single-scan evaluation. The vertical FOV is $[-25^{\circ},3^{\circ}]$. nuScenes [47] is a multimodal dataset that is widely used in autonomous driving scenarios. It contains LiDAR point clouds captured around streets in Boston and Singapore, provided by 32-beam LiDAR sensors. The dataset includes $1,000$ driving scenes with sparser points. It is annotated with $32$ classes and only $16$ semantic categories are used for evaluation. The vertical FOV is $[-30^{\circ},10^{\circ}]$. SemanticPOSS [68] is a more challenging benchmark collected at Peking University and features much smaller and sparser point clouds. It consists of $2,988$ LiDAR scenes with numerous dynamic instances. Officially, it is divided into $6$ sequences, with sequence $2$ used for evaluation with $13$ merged categories. The vertical FOV is $[-16^{\circ},7^{\circ}]$. ScribbleKITTI [67] shares the same scenes with SemanticKITTI [32] but only provides weak annotations with line scribbles. It contains $189$ million labeled points, with approximately $8.06\%$ labeled points available during training. Additionally, we conduct a comprehensive out-of-distribution robustness evaluation experiment on SemanticKITTI-C and nuScenes-C from the Robo3D [76] benchmark. Each dataset contains eight corruption types, including “fog”, “wet ground”, “snow”, “motion blur”, “beam missing”, “crosstalk”, “incomplete echo”, and “cross sensor”.
@@
Implementation Details. We implement FRNet using the popular MMDetection3D [77] codebase and employ CENet [61] as the 2D backbone to extract frustum features. For SemanticKITTI [32] and ScribbelKITTI [67], the point cloud is divided into $64\times 512$ frustum regions, while for nuScenes [47] and SemanticPOSS [68], it is divided into $32\times 480$. For optimization, we choose AdamW [78] as our default optimizer with an initial learning rate of $0.01$. We utilized the OneCycle scheduler [79] to dynamically adjust the learning rate during training. All models are trained using four GPUs for $50$ epochs on SemanticKITTI [32] and ScribbelKITTI [67], and for $80$ epochs on nuScenes [47] and SemanticPOSS [68], respectively. The batch size is set to $4$ for each GPU. Additionally, we implement a faster variant, dubbed Fast-FRNet, which features a larger frustum area and a scaled-down 2D backbone. Specifically, the point cloud is divided into $32\times 360$ frustum regions for all datasets, and the 2D backbone is a variant of ResNet-18 [80], following the design of CENet [61].
@@ IV-B Comparative Study
Comparisons to State of the Arts. We begin by comparing FRNet with state-of-the-art LiDAR segmentation methods across different representations, including sparse-voxel, point-view, multi-view, and range-view representations, on both fully-supervised and weakly-supervised benchmarks. Tab. I provides a summary of all the results. In this benchmark, both FrustumMix and RangeInterpolation are employed as data augmentation techniques to enhance model performance. To ensure a fair comparison, we evaluate the Frame Per Second (FPS) on a single GeForce RTX 2080Ti GPU for all models. For SemanticKITTI [32] and nuScenes [47], we report mIoU scores on both the validation and test sets. For ScribbleKITTI [67] and SemanticPOSS [68], which share the same data across the val and test sets, we report both mIoU and mAcc scores on the val set. Compared with recent multi-view methods [24], although FRNet does not exhibit a significant advantage in performance, it shows notable superiority in terms of parameters and FPS. In comparison with range-view, sparse-voxel, and raw-points methods, FRNet achieves appealing performance while still maintaining high efficiency. Additionally, Fast-FRNet demonstrates the fastest inference speed with only a minor drop in performance.
@@
Label-Efficient LiDAR Semantic Segmentation. Recently, semi-supervised learning has gained traction in LiDAR segmentation. In this work, we adopt the approach of LaserMix [48] to apply FRNet to the semi-supervised segmentation task, as shown in Tab. II. For fair comparisons with other methods, FrustumMix and RangeInterpolation are not employed as data augmentation techniques in this benchmark, and TTA is excluded during inference. We first utilize Cylinder3D [16] and PTv3 [12] as backbones and leverage the proposed FrustumMix technique to mix labeled and unlabeled data for consistency learning, which aligns with the LaserMix framework [48]. Notably, FrustumMix demonstrates superior performance compared to the original LaserMix [48] strategy. Next, we employ FRNet as the backbone and evaluate various strategies for mixing labeled and unlabeled scans, including PolarMix [49], LaserMix [48], and our FrustumMix. The Sup.-only baseline results are also obtained without FrustumMix and RangeInterpolation as data augmentation techniques. The results indicate that FrustumMix consistently achieves the best scores across all data splits when compared with other mixing strategies, particularly with limited annotations.
Out-of-Distribution Robustness. Robustness, which reflects a model’s ability to generalize under different conditions, is crucial for automotive security systems, particularly in extreme weather conditions [81, 82]. To evaluate the out-of-distribution robustness of FRNet, we utilize the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [76]. In this benchmark, the models trained on SemanticKITTI [32] and nuScenes [47] are directly applied to evaluate the performance on SemanticKITTI-C and nuScenes-C [76], respectively, without any fine-tunning. As shown in Tab. III, FRNet achieves promising performance across most corruption types and demonstrates superiority over recent works employing various LiDAR representations, including sparse voxel [17, 15, 16], range view [59, 61], multi-view fusion [71], and raw points [11].
@@
IV-C Ablation Study
TABLE IV: Ablation study of each component in FRNet on the val set of SemanticKITTI [32] and nuScenes [47]. FFE: Frustum Feature Encoder. FP: Frustum-Point Fusion. FH: Fusion Head. FS: Frustum-level Supervision. RI: RangeInterpolation. TTA: Test Time Augmentation. All mIoU and mAcc scores are given in percentage (%).
@@ IV-C Ablation Study
Component Designs. Tab. IV summarizes the ablation results of each component in the FRNet architecture. Notably, FrustumMix is employed across all setups in this ablation study. Firstly, we directly apply the frustum feature encoder to CENet [61], incorporating KNN post-processing to predict semantic labels over the entire point cloud, achieving $62.7\%$ and $72.8\%$ mIoU on SemanticKITTI [32] and nuScenes [47], respectively. Next, by integrating point-frustum fusion to hierarchically update point features, the semantic results are predicted with the updated point features in an end-to-end manner, resulting in improvements of $1.6\%$ and $0.8\%$ mIoU, respectively. To regularize the frustum features, the frustum-level supervision is introduced and it leads to significant gains of $1.9\%$ and $2.8\%$ mIoU, respectively. The fusion head module, which fuses multiple features at different levels for more accurate prediction, brings an additional $0.8\%$ mIoU gain on both datasets. Furthermore, the proposed RangeInterpolation, aimed at alleviating empty pixels in the 2D representation by reconstructing semantic surfaces via surrounding range information, provides a performance boost of approximately $0.5\%$ mIoU. Finally, adopting Test Time Augmentation during inference, as demonstrated in prior works, leads to improvements of $1.1\%$ and $1.3\%$ mIoU, respectively.
[…]
Number of Frustums Height $(H)$ Width $(W)$ mIoU mAcc
$131,072$ $64$ $2048$ $66.8$ $73.8$
$65,536$ $64$ $1024$ $67.1$ $73.4$
$65,536$ $32$ $2048$ $65.5$ $72.7$
$32,768$ $64$ $512$ $\mathbf{67.6}$ $\mathbf{74.2}$
$32,768$ $32$ $1024$ $66.0$ $72.4$
$16,384$ $64$ $256$ $64.3$ $70.5$
$16,384$ $32$ $512$ $65.5$ $72.1$
$8,192$ $32$ $256$ $64.3$ $70.8$
Frustum Representation Resolution. Finding a suitable frustum resolution is crucial for balancing accuracy and efficiency. In Tab. V, we conduct a series of experiments with various frustum resolutions. While higher resolution can capture more detailed features, it may not always lead to optimal results. The limited area of the frustum region can result in a lack of meaningful structure, causing the model to struggle with noisy or sparse data. Conversely, although lower resolution can improve inference efficiency and provide a more coherent representation, the larger area of the frustum region may introduce excessive noise information, particularly at object boundaries. Thus, finding an optimal frustum resolution is crucial for balancing the richness of features with computational demands and maintaining overall model performance.
@@
TABLE VI: Ablation study on window settings in RangeInterpolation on the val set of SemanticKITTI [32]. Green pixel denotes the empty pixel needed to be interpolated. Violet pixels are the surrounding pixels used to generate the point. All mIoU scores are given in percentage (%).
@@
RangeInterpolation Window Settings. RangeInterpolation reconstructs semantic surfaces based on the range image, facilitating a compact 2D representation for coherent semantic learning. Tab. VI showcases the performance of RangeInterpolation with different window settings. Notably, we observe consistent performance improvements when considering interpolation along the horizontal or vertical direction exclusively. Specifically, configurations employing horizontal windows, such as the $1\times 3$ and $1\times 5$ settings, demonstrate superior results compared to those focusing on the vertical direction ($3\times 1$ and $5\times 1$ windows). We conjecture that points aligned along the same laser beams often share similar attributes, leading to more accurate and coherent semantic surfaces. Conversely, larger window sizes, such as the $3\times 5$, $3\times 3$, and $5\times 5$ configurations, tend to exhibit non-uniform point distributions within the window, resulting in noisy points that fail to accurately represent the expected region and consequently leading to a drop in performance.
@@
TABLE VII: Inference efficiency comparison of FRNet and MinkUNet [17] implemented with various codebases and sparse convolution backends, as well as PTv3 [12] on the val set of SemanticKITTI [32].
Codebase Backbone Backend FPS
MMDetection3D MinkUNet MinkowskiEngine $9.1$
MinkUNet SpConv $8.6$
MinkUNet TorchSparse $10.2$
MinkUNet TorchSparse++ $21.3$
FRNet Convolution $29.1$
Fast-FRNet Convolution $\mathbf{33.8}$
PCSeg MinkUNet TorchSparse $10.6$
PointCept MinkUNet MinkowskiEngine $8.9$
PTv3 FlashAttention $20.2$
Inference Efficiency. Multiple sparse convolution backends, including MinknowskiEngine, SpConv, TorchSparse, and TorchSparse++, offer efficient processing for voxel-based methods. To compare FRNet with these backends, we implement MinkUNet [17] using the MMDetection3D codebase across various backends. Additionally, we evaluate the FPS of MinkUNet implemented with the PCSeg and PointCept codebases, as well as the recent efficient method PTv3 [12]. All evaluations are conducted on a single NVIDIA 2080Ti GPU. As shown in Tab. VII, while TorchSparse++ delivers satisfactory efficiency for LiDAR segmentation, FRNet achieves the fastest inference speed, making it suitable for real-time applications in autonomous driving systems.
(a) car
(b) person
(c) vegetation
(d) drivable surface
(e) sidewalk
(f) trailer
Figure 10: The open-vocabulary LiDAR semantic segmentation experiments of FRNet. Our model demonstrates the ability to perform open-vocabulary segmentation with various textual inputs. The segmented points corresponding to the provided text descriptions are highlighted in blue, showcasing its potential for open-world perception tasks.
TABLE VIII: The multi-source training results of the proposed Fast-FRNet on the val set of the SemanticKITTI [32], nuScenes [47], and SemanticPOSS [68] datasets. All mIoU and mAcc scores are given in percentage (%).
Method SemKITTI nuScenes SemPOSS
mIoU mAcc mIoU mAcc mIoU mAcc
Single-Source $67.1$ $73.8$ $78.8$ $84.4$ $52.5$ $67.1$
Multi-Source $67.4$ $74.2$ $79.1$ $84.7$ $53.2$ $68.0$
Fast-FRNet (w/ PPT [83]) $\mathbf{68.2}$ $\mathbf{74.7}$ $\mathbf{79.3}$ $\mathbf{84.9}$ $\mathbf{53.9}$ $\mathbf{68.4}$
TABLE IX: Zero-shot segmentation performance on the val set of nuScenes [47]. All mIoU and mAcc scores are given in percentage (%).
Method nuScenes
mIoU mAcc
OpenScene [84] $42.1$ $61.8$
+ FRNet $\mathbf{43.5}$ $\mathbf{62.0}$
Multi-Dataset Fusion. Recent studies [83, 51] highlight the importance of large-scale benchmarks achieved through multi-dataset fusion strategies across heterogeneous datasets. In this work, We extend Fast-FRNet to support multi-dataset joint training, as shown in Tab. VIII. Initially, we combine multi-source datasets using mean-variance normalization to align intensity values, which provides slight performance gains compared to single-source training. Additionally, inspired by PPT [83], we incorporate point-prompt techniques to facilitate multi-source joint training. This approach yields further improvements, demonstrating the effectiveness of prompt-guided strategies in multi-dataset fusion.
Open-Vocabulary Semantic Segmentation. To investigate the potential of FRNet in open-vocabulary segmentation, we adopt an approach inspired by OpenScene [84]. Specifically, we integrate text and image features into the framework, aligning them with point features. During inference, text inputs are used to guide the segmentation of point clouds with similar features. As shown in Tab. IX, FRNet demonstrates promising performance in zero-shot segmentation. Additionally, in Fig. 10, we present qualitative results of segmentation using corresponding text inputs, highlighting the capability of FRNet to handle open-world perception tasks effectively.